OpenStudy (anonymous):

Can someone Explain this to me please. 5x-3y-16=0 4x-5y-5=0 Solution Set x| 2 |5 5(2)-36-16 5(5)-3y-16=0 y|-2 |3 10-3y-16 25-36-16=0 3y=16-10 36=16-25 3y=6 3y=9 3 3 y=-2 3 3 y=3 x|5|0 4(5)-5y-5=0 -5y-5=0 y|3|-1 20-5y=0 -5y=5

OpenStudy (anonymous):

the same as 5x-3y=16 4x-5y=5 now to eliminate y we multiply the first equation by 5 and then the second with 3 and get 25x-15y=80 12x-15y=15 now subtract the first minus the second (25x-12x)+(15y-(-15y))=80-15 13x+0=65 13x=65 x=5 now from 5x-3y=16 5(5)-3y=16 25-16=3y 9=3y 3=y